OpenStudy (anonymous):

Can you help me solve this? -9(a + 16) + -4 = 23

OpenStudy (anonymous):

Distribute -9 to the parenthesis first..

OpenStudy (anonymous):

-9a-144=23+4 -9a=27+144 -9a=331 (divide by -9) a=-19 .